What is the energy of a photon of light of wavelength 0.90 µm?
Answer #1

Solution)

Given,

Wavelength, lambda=0.9*10^-6 m

We know, Energy, E= hc/lambda

Substitute values,

E=(6.626*10^-34)*(3*10^8)/(0.9*10^-6)

E= 2.20*10^-19 J (Ans)

==============